Michael and Anna intend to buy a birthday present for Carol. They decide that the cost of the present should not be more than $20 and than Anna will pay $2 more than Michael. What is the maximum amount that Elvin has to pay?
Answer by josgarithmetic(39623)

How Much these two pay.
n Anna
m Michael

m+n=20 and n-m=2; no pay data was given about Elvin. This person is not part of the problem description.

n=m+2;
m+m+2=20
2m+2=20
m+1=10
m=9 and n=11.

Michael pays $9 and Anna pays $11.
The quantity to pay is accounted according to the description. Elvin pays nothing.